# Pralekha: Cross-Lingual Document Alignment for Indic Languages

### v2

@@ Abstract
Mining parallel document pairs for document-level machine translation (MT) remains challenging due to the limitations of existing Cross-Lingual Document Alignment (CLDA) techniques. Most approaches rely on metadata such as URLs, which is often unavailable in low-resource language settings, while others represent documents using pooled sentence embeddings, which fail to capture fine-grained alignment cues. Moreover, current sentence embedding models have limited context windows, hindering their ability to represent document-level information effectively. To address these challenges for Indic languages, we introduce Pralekha [^1] Pralekha means document in Sanskrit, an ancient Indo-Aryan language spoken in India., a large-scale benchmark for evaluating document-level alignment techniques. It contains over 3 million aligned document pairs across 11 Indic languages and English, of which 1.5 million are English–Indic pairs. Furthermore, we propose Document Alignment Coefficient (DAC), a novel metric for fine-grained document alignment. Unlike pooling-based approaches, DAC aligns documents by matching smaller chunks and computes similarity as the ratio of aligned chunks to the average number of chunks in a pair. Intrinsic evaluation shows that DAC achieves substantial improvements over pooling-based baselines, particularly in noisy scenarios. Extrinsic evaluation further demonstrates that document MT models trained on DAC-aligned pairs consistently outperform those using baseline alignment methods. These results highlight DAC’s effectiveness for parallel document mining. The Pralekha dataset and CLDA evaluation framework will be made publicly available.
@@ 1 Introduction
The advent of Large Language Models (LLMs) has significantly improved the ability to process long-context information, enabling advancements in document-level tasks such as machine translation (MT). Document-level MT extends beyond sentence-level translation by capturing discourse-level dependencies, coherence, and contextual information across entire documents, leading to more accurate and fluent translations. While there are efforts Ramesh et al. (2022); Siripragada et al. (2020) that provide sentence-level parallel data, large-scale document-level parallel corpora remain scarce Sannigrahi et al. (2023); Pal et al. (2024).
@@
To address this gap, parallel document mining is essential for building high-quality training corpora for document-level MT. However, this progress is hindered by two key challenges: (a) the lack of large-scale document alignment evaluation benchmarks and (b) the limitations of current Cross-Lingual Document Alignment (CLDA) techniques. Reliable evaluation benchmarks are crucial because the effectiveness of an alignment method cannot be improved without accurate performance measurements. As for the limitations of CLDA, most existing embedding models were developed for sentence-level parallel corpus mining Schwenk et al. (2020, 2021); Bañón et al. (2020). As a result, they have limited context windows and can represent only parts of a document Zhu et al. (2024). Document-level embeddings are therefore often obtained by truncating documents or pooling embeddings of independently encoded chunks, both of which can be lossy Sannigrahi et al. (2023). These approaches overlook chunk-level cues, which reduces their ability to capture complete document semantics.
@@
In this paper, we focus on Indic languages, which have abundant sentence-level parallel corpora and benchmarks  Ramesh et al. (2022); Siripragada et al. (2020); Gala et al. (2023) but remain low-resource for CLDA and Document-level MT. We first introduce Pralekha, a large-scale parallel document dataset for evaluating CLDA techniques across 11 Indic languages and English. Pralekha comprises over 3 million document pairs between 11 Indic languages and English, of which 1.5 million are English–Indic pairs. These are high-quality, human-verified aligned documents sourced from reliable domains such as the Indian Press Information Bureau (PIB) and the Mann Ki Baat (MKB) radio program. It serves both as a benchmark for assessing CLDA and as a domain-specific parallel corpus for training document-level MT models for Indic languages. Alongside Pralekha, we propose Document Alignment Coefficient (DAC), a novel metric for CLDA at the fine-granular level. Unlike existing approaches that rely on pooled document-level embeddings, we focus on smaller granular units, called chunks. We first embed and align these chunks, then compute the DAC of a document pair as the ratio of aligned chunks to the average number of chunks in the pair.
@@
In our experiments using Pralekha under realistic noisy document mining settings, we evaluate the proposed DAC approach against existing pooling-based baselines. DAC achieves average precision, recall, and F1 scores of 0.8932, 0.6312, and 0.7372, respectively, showing improvements of 15–20% in precision and 5–10% in F1 score over the baselines. We further benchmark DAC on the CCAligned El-Kishky et al. (2020) dataset, where it attains average precision, recall, and F1 scores of 0.8110, 0.6511, and 0.7203, respectively, outperforming the baselines across all metrics. Having established DAC’s intrinsic effectiveness, we conduct an extrinsic evaluation by training document-level MT models on mined documents using open-source LLMs. The results show that MT models trained on DAC-aligned pairs consistently outperform those trained with baseline alignment methods. The Pralekha dataset and CLDA evaluation framework will be made publicly available.
@@ 2 Related Work
Cross-Lingual Sentence Alignment. Early approaches to sentence-level alignment, such as Europarl Koehn (2005), relied on manually curated metadata Abdul-Rauf and Schwenk (2009); Do et al. (2009). Recent advancements in multilingual embeddings have significantly improved large-scale mining of parallel data from unstructured sources. Methods such as LaBSE Feng et al. (2022), LASER Schwenk and Douze (2017), SONAR Duquenne et al. (2023), and hierarchical bilingual retrieval Guo et al. (2019) leverage multilingual sentence representations to enhance cross-lingual alignment accuracy. These techniques have enabled large-scale efforts such as CCMatrix Schwenk et al. (2020), WikiMatrix Schwenk et al. (2021), ParaCrawl Bañón et al. (2020), NLLB Team et al. (2022), and BPCC Gala et al. (2023), which extract and align sentence pairs from massive web crawls.
Cross-Lingual Document Alignment. For document-level alignment, early approaches relied on metadata rather than content, using features such as URL and page structure similarity Resnik and Smith (2003). While multilingual encoders have improved sentence-level parallel data extraction, document-level alignment remains challenging due to the limited context windows of existing sentence embedding models Zhu et al. (2024). In this work, we explore a chunk-based method that optimally leverages fine-granular embeddings for document-level alignment.
Indic Language Corpora. Despite India’s large population (1.43 billion), its languages remain low-resource, particularly for high-quality document-level data. Significant efforts have been made to build monolingual corpora, such as IndicNLP Suite Kakwani et al. (2020) and Sangraha Khan et al. (2024), as well as parallel corpora and benchmarks like IN22 and BPCC Gala et al. (2023). However, no publicly available cross-lingual parallel document datasets exist for Indic languages Sannigrahi et al. (2023). The only work referencing parallel documents uses them solely for mining sentence pairs, without releasing the document-level data Siripragada et al. (2020). This underscores the urgent need to construct parallel document corpora for cross-lingual document alignment and document-level machine translation, a gap that our work aims to fill.
@@ 3 Pralekha
Figure 3 presents statistics on the average number of sentences per document for each language in Pralekha. The average sentence count ranges between 20 and 30, indicating that the documents are relatively long, making our benchmark challenging for CLDA.
Pralekha covers two broad domains - News Bulletins and Podcast Scripts, therefore containing a mixture of written as well as spoken form of data. Following the methodology of Sangraha Khan et al. (2024), we custom-scraped data from the Indian Press Information Bureau (PIB) [^2] https://pib.gov.in website, aligning documents by matching bulletin IDs interlinking bulletins across languages. These documents are manually written and hence are of high quality. For podcast scripts, we employed the approach used by Khan et al. (2024) and Siripragada et al. (2020) to collect transcripts from Mann Ki Baat, [^3] https://www.pmindia.gov.in/en/mann-ki-baat a radio program hosted by the Indian Prime Minister. This program is typically spoken in Hindi, then manually transcribed and translated into various other Indian languages. Sometimes, the website’s meta-data incorrectly marks the document’s languages, which we fix with a simple combination of IndicLID Madhani et al. (2023) and script-specific Unicode ranges.
@@ 4 Cross-Lingual Document Alignment
We next describe the evaluation framework used to benchmark various CLDA techniques. As illustrated in Figure 1, the framework includes our proposed approach, which leverages the Document Alignment Coefficient (DAC), along with the pooling-based baseline methods. The following subsections discuss each component of the framework in detail.
@@ 4.1 Granularity
The granularity level $G$ defines the unit of text at which alignment is performed and is a key parameter in our study. We evaluate document alignment at three levels of granularity: $G=1$ (sentence-level, as in Schwenk et al. (2020)), $G=2,4,8$ (chunk-level, where each chunk consists of 2, 4, or 8 sentences, respectively), and $G=|D|$ (document-level, as in El-Kishky et al. (2020); El-Kishky and Guzmán (2020)).
At one extreme, $G=1$ produces too many embeddings to align. This makes the process inefficient and provides limited context for semantic comparison. At the other extreme, $G=|D|$ leads to overly coarse representations and information loss. It also creates challenges in computing a single document-level embedding because embedding models have a limited context window Zhu et al. (2024). Both extremes have limitations; therefore, we consider intermediate values of $G$ that strike a balance between efficiency and information retention. A larger granularity helps preserve local coherence and capture broader document-level information, addressing the shortcomings of traditional sentence-based approaches Xie et al. (2024).
Pooling Strategies for Baselines. We benchmark several pooling strategies used in prior work El-Kishky et al. (2020); El-Kishky and Guzmán (2020) to construct document embeddings ($G=|D|$) by aggregating finer-grained unit embeddings.
@@ 4.2 Embedding Model
To compute multilingual embeddings we consider LaBSE [Uncaptioned image] Feng et al. (2022) and SONAR [Uncaptioned image] Duquenne et al. (2023). LaBSE is designed for cross-lingual sentence retrieval, leveraging both parallel and monolingual data to generate language-agnostic embeddings. SONAR extends this approach by training on a more diverse set of languages, leading to improved representation quality for both high-resource and low-resource languages. Computing Embeddings. After segmenting each document into units based on the chosen granularity $G$, we obtain representations for each segment by encoding the entire unit with the embedding model. Because the values of $G$ considered in our study are relatively small, each segment fits within the context window of the embedding models used.
@@ 4.3 Alignment Algorithm
[…]
@@
[…]
@@ 4.4 Document Alignment Coefficient
The DAC produces a normalized alignment score between 0 and 1, where higher values indicate stronger alignment, and a threshold on this score can be set to control the final alignment quality. A lower DAC threshold allows a balance between quality and quantity, accepting some misalignments while increasing coverage. On the other hand, a higher threshold prioritizes precision, ensuring fewer incorrect alignments at the cost of reduced yield. The threshold can be adjusted based on the specific requirements of the aligned dataset and its intended downstream task. A detailed analysis of various DAC Thresholds is provided in Figure 7 and Table 3 in Appendix A
@@ 5.1 Intrinsic Evaluation
Pralekha contains only parallel (alignable) documents, so aligning only these would not reflect a realistic and challenging scenario. To simulate real-world multilingual corpora setting, we sample random (unalignable) documents from Sangraha Unverified Khan et al. (2024) that cannot be aligned with those in Pralekha and inject them as noise into the dataset. We select a number of unalignable documents equal to 50% of the documents in Pralekha, resulting in a 1:2 ratio of unalignable to alignable documents during alignment. Table 1 reports the counts of unique alignable and unalignable documents for each language used in our experiments. Additionally, we compare the best-performing baseline with our proposed DAC approach on the CCAligned dataset El-Kishky et al. (2020), using a similar data setting.
@@
[…]
@@
To assess the intrinsic performance of various document alignment methods, we evaluate three key metrics: precision, recall, and F1 score. While a high F1 score is generally desirable as it balances precision and recall, the ideal metric depends on the specific downstream task.
5.2 Extrinsic Evaluation
Data Setting :
Implementation and Training :
We fine-tune two open-source LLMs: Llama-3.2-1B666huggingface.co/meta-llama/Llama-3.2-1B [Uncaptioned image] and Sarvam-1777huggingface.co/sarvamai/sarvam-1 [Uncaptioned image]. LLaMA-3.2-1B is a 1 billion parameter multilingual model released by Meta AI, while Sarvam-1 is a 2 billion parameter model developed by Sarvam AI, optimized for Indic languages. We used open-instruct [^8] https://github.com/allenai/open-instruct for full fine-tuning using the prompt-completion template. We used default hyperparameters recommended in open-instruct, most notably: learning rates of $5\times 10^{-5}$, maximum context (prompt+completion) lengths of 4096, Adam optimizer and no weight decay. We trained only bilingual models, a single model for English $\leftrightarrow$ X translation, for a total of 16 models (8 languages, 2 alignment approaches). Our models were trained for 1 epoch. We did greedy decoding with a maximum of 4,096 generated tokens.

### v3

@@ Abstract
Mining parallel document pairs for document-level machine translation (MT) remains challenging due to the limitations of existing Cross-Lingual Document Alignment (CLDA) techniques. Existing methods often rely on metadata such as URLs, which are scarce, or on pooled document representations that fail to capture fine-grained alignment cues. Moreover, the limited context window of sentence embedding models hinders their ability to represent document-level context, while sentence-based alignment introduces a combinatorially large search space, leading to high computational cost. To address these challenges for Indic languages, we introduce Pralekha [^1] Pralekha means document in Sanskrit, an ancient Indo-Aryan language., a benchmark containing over 3 million aligned document pairs across 11 Indic languages and English, which includes 1.5 million English–Indic pairs. Furthermore, we propose Document Alignment Coefficient (DAC), a novel metric for fine-grained document alignment. Unlike pooling-based methods, DAC aligns documents by matching smaller chunks and computes similarity as the ratio of aligned chunks to the average number of chunks in a pair. Intrinsic evaluation shows that our chunk-based method is 2–3× faster while maintaining competitive performance, and that DAC achieves substantial gains over pooling-based baselines. Extrinsic evaluation further demonstrates that document-level MT models trained on DAC-aligned pairs consistently outperform those using baseline alignment methods. These results highlight DAC’s effectiveness for parallel document mining. The dataset and evaluation framework are publicly available to support further research.
@@ 1 Introduction
The advent of Large Language Models (LLMs) has significantly improved the ability to process long-context information, enabling advancements in document-level tasks such as machine translation (MT). Document-level MT extends beyond sentence-level translation by capturing discourse-level dependencies, coherence, and contextual information across entire documents, leading to more accurate and fluent translations. While there are efforts ramesh-etal-2022-samanantar; siripragada-etal-2020-multilingual that provide sentence-level parallel data, large-scale document-level parallel corpora remain scarce sannigrahi-etal-2023-best; pal-etal-2024-document.
@@
To address this gap, parallel document mining is essential for building high-quality training corpora for document-level MT. However, this progress is hindered by two key challenges: (a) the lack of large-scale document alignment evaluation benchmarks and (b) the limitations of current Cross-Lingual Document Alignment (CLDA) techniques. Reliable evaluation benchmarks are crucial because the effectiveness of an alignment method cannot be improved without accurate performance measurements. As for the limitations of CLDA, most existing embedding models were developed for sentence-level parallel corpus mining schwenk2020ccmatrixminingbillionshighquality; schwenk-etal-2021-wikimatrix; banon-etal-2020-paracrawl. As a result, they have limited context windows and can represent only parts of a document zhu2024longembedextendingembeddingmodels. Document-level embeddings are therefore often obtained by truncating documents or pooling embeddings of independently encoded chunks, both of which can be lossy sannigrahi-etal-2023-best. These approaches overlook chunk-level cues, which reduces their ability to capture complete document semantics. Even without pooling, the reliance of existing alignment techniques on sentence-level operations leads to combinatorial growth in the search space, making large-scale document alignment computationally expensive.
@@
In this paper, we focus on Indic languages, which have abundant sentence-level parallel corpora and benchmarks  ramesh-etal-2022-samanantar; siripragada-etal-2020-multilingual; gala2023indictrans2highqualityaccessiblemachine but remain low-resource for CLDA and Document-level MT. We first introduce Pralekha, a large-scale parallel document dataset for evaluating CLDA techniques across 11 Indic languages and English. Pralekha comprises over 3 million document pairs between 11 Indic languages and English, of which 1.5 million are English–Indic pairs. These are high-quality, human-verified aligned documents sourced from reliable domains such as the Indian Press Information Bureau (PIB) and the Mann Ki Baat (MKB) radio program. It serves both as a benchmark for assessing CLDA and as a domain-specific parallel corpus for training document-level MT models for Indic languages. Alongside Pralekha, we propose Document Alignment Coefficient (DAC), a novel metric for CLDA at the fine-granular level. Unlike existing approaches that rely on pooled document-level embeddings, we focus on smaller granular units, called chunks. We first embed and align these chunks, then compute the DAC of a document pair as the ratio of aligned chunks to the average number of chunks in the pair.
@@
In our experiments using Pralekha under realistic noisy document mining settings, we evaluate the proposed DAC approach against existing pooling-based baselines. DAC achieves average precision, recall, and F1 scores of 0.8932, 0.6312, and 0.7372, respectively, showing improvements of 15–20% in precision and 5–10% in F1 score over the baselines. We further benchmark DAC on the CCAligned el-kishky-etal-2020-ccaligned dataset, where it attains average precision, recall, and F1 scores of 0.8110, 0.6511, and 0.7203, respectively, outperforming the baselines across all metrics. Additionally, our chunk-based method demonstrates a 2–3× speedup over sentence-based alignment techniques while maintaining competitive performance. Having established DAC’s intrinsic effectiveness, we conduct an extrinsic evaluation by training document-level MT models on mined documents using open-source LLMs. The results show that MT models trained on DAC-aligned pairs consistently outperform those trained with baseline alignment methods. The Pralekha dataset and CLDA evaluation framework is publicly available to support further research in this area.
@@ 2 Related Work
Cross-Lingual Sentence Alignment. Early approaches to sentence-level alignment, such as Europarl koehn-2005-europarl, relied on manually curated metadata abdul-rauf-schwenk-2009-improving; do-etal-2009-mining. Recent advancements in multilingual embeddings have significantly improved large-scale mining of parallel data from unstructured sources. Methods such as LaBSE feng-etal-2022-language, LASER schwenk-douze-2017-learning, SONAR duquenne2023sonarsentencelevelmultimodallanguageagnostic, and hierarchical bilingual retrieval guo-etal-2019-hierarchical leverage multilingual sentence representations to enhance cross-lingual alignment accuracy. These techniques have enabled large-scale efforts such as CCMatrix schwenk2020ccmatrixminingbillionshighquality, WikiMatrix schwenk-etal-2021-wikimatrix, ParaCrawl banon-etal-2020-paracrawl, NLLB nllbteam2022languageleftbehindscaling, and BPCC gala2023indictrans2highqualityaccessiblemachine, which extract and align sentence pairs from massive web crawls.
Cross-Lingual Document Alignment. For document-level alignment, early approaches relied on metadata rather than content, using features such as URL and page structure similarity resnik-smith-2003-web. While multilingual encoders have improved sentence-level parallel data extraction, document-level alignment remains challenging due to the limited context windows of existing sentence embedding models zhu2024longembedextendingembeddingmodels. In this work, we explore a faster chunk-based method that effectively leverages fine-granular embeddings for document-level alignment.
Indic Language Corpora. Despite India’s large population (1.43 billion), its languages remain low-resource, particularly for high-quality document-level data. Significant efforts have been made to build monolingual corpora, such as IndicNLP Suite kakwani-etal-2020-indicnlpsuite and Sangraha khan-etal-2024-indicllmsuite, as well as parallel corpora and benchmarks like IN22 and BPCC gala2023indictrans2highqualityaccessiblemachine. However, no publicly available cross-lingual parallel document datasets exist for Indic languages sannigrahi-etal-2023-best. The only work referencing parallel documents uses them solely for mining sentence pairs, without releasing the document-level data siripragada-etal-2020-multilingual. This underscores the urgent need to construct parallel document corpora for cross-lingual document alignment and document-level machine translation, a gap that our work aims to fill.
@@ 3 Pralekha
Figure 3 presents statistics on the average number of sentences per document for each language in Pralekha. The average sentence count ranges from 20 to 30, indicating that the documents are relatively long and making the benchmark more challenging.
Pralekha covers two broad domains - News Bulletins and Podcast Scripts, therefore containing a mixture of written as well as spoken form of data. Following the methodology of Sangraha khan-etal-2024-indicllmsuite, we custom-scraped data from the Indian Press Information Bureau (PIB) [^2] https://pib.gov.in website, aligning documents by matching bulletin IDs interlinking bulletins across languages. These documents are manually written and hence are of high quality. For podcast scripts, we employed the approach used by khan-etal-2024-indicllmsuite and siripragada-etal-2020-multilingual to collect transcripts from Mann Ki Baat, [^3] https://www.pmindia.gov.in/en/mann-ki-baat a radio program hosted by the Indian Prime Minister. This program is typically spoken in Hindi, then manually transcribed and translated into various other Indian languages. Sometimes, the website’s meta-data incorrectly marks the document’s languages, which we fix with a simple combination of IndicLID madhani-etal-2023-bhasa-abhijnaanam and script-specific Unicode ranges.
@@ 4 Cross-Lingual Document Alignment
We next describe the evaluation framework used to benchmark various CLDA techniques. As illustrated in Figure 1, the framework includes our proposed approach, which leverages the Document Alignment Coefficient (DAC), along with the pooling-based baseline methods. The following subsections discuss each component of the CLDA evaluation framework in detail.
@@ 4.1 Granularity
The granularity level $G$ defines the unit of text at which alignment is performed and is a key parameter in our study. We evaluate document alignment at three levels of granularity: $G=1$ (sentence-level, as in schwenk2020ccmatrixminingbillionshighquality), $G=2,4,8$ (chunk-level, where each chunk consists of 2, 4, or 8 sentences, respectively), and $G=|D|$ (document-level, as in el-kishky-etal-2020-ccaligned; el-kishky-guzman-2020-massively).
At one extreme, $G=1$ produces too many embeddings to align, resulting in a combinatorially large search space that leads to high computational cost, while offering limited context for semantic comparison. At the other extreme, $G=|D|$ leads to overly coarse representations and information loss. It also creates challenges in computing a single document-level embedding because embedding models have a limited context window zhu2024longembedextendingembeddingmodels. Both extremes have limitations; therefore, we consider intermediate values of $G$ that strike a balance between efficiency and information retention. A larger granularity helps preserve local coherence and capture broader document-level information, addressing the shortcomings of traditional sentence-based approaches xie2024chunkalignselectsimple.
Pooling Strategies for Baselines. We benchmark several pooling strategies used in prior work el-kishky-etal-2020-ccaligned; el-kishky-guzman-2020-massively to construct document embeddings ($G=|D|$) by aggregating finer-grained unit embeddings.
@@ 4.2 Embedding Model
To compute multilingual embeddings we consider LaBSE [Uncaptioned image] feng-etal-2022-language and SONAR [Uncaptioned image] duquenne2023sonarsentencelevelmultimodallanguageagnostic. LaBSE is designed for cross-lingual sentence retrieval, leveraging both parallel and monolingual data to generate language-agnostic embeddings. SONAR extends this approach by training on a more diverse set of languages, leading to improved representation quality for both high-resource and low-resource languages. Computing Embeddings. After segmenting each document into units based on the chosen granularity $G$, we obtain representations for each segment by encoding the entire unit with the embedding model. Because the values of $G$ considered in our study are relatively small, each segment fits within the context window of the embedding models used.
@@ 4.3 Alignment Algorithm
After computing unit embeddings at various text granularities from the monolingual corpora in the source and target languages, we apply the Alignment Algorithm 1 to align these embeddings. The algorithm performs margin-based bitext mining in a shared multilingual space to extract parallel units schwenk2020ccmatrixminingbillionshighquality. It adopts the max-margin criterion, which has been shown to improve alignment quality compared to an absolute similarity threshold artetxe-schwenk-2019-margin. To build efficient similarity indices, we use the FAISS library [^4] https://github.com/facebookresearch/faiss. For each unit embedding, we retrieve the top-$k$ nearest neighbors and compute margin scores $M(x,y)$ to capture similarity in a relative context. Candidate pairs are then ranked by their margin scores, and a greedy bipartite matching strategy is applied to select the highest-scoring pairs, ensuring that each embedding is aligned only once.
@@
$\texttt{index\_X}\leftarrow\texttt{IndexFlatIP}(\mathbf{X})$
$\texttt{index\_Y}\leftarrow\texttt{IndexFlatIP}(\mathbf{Y})$
4:Query top-$k$ nearest neighbors:
@@ 4.4 Document Alignment Coefficient
The DAC produces a normalized alignment score between 0 and 1, where higher values indicate stronger alignment, and a threshold on this score can be set to control the final alignment quality. A lower DAC threshold allows a balance between quality and quantity, accepting some misalignments while increasing coverage. On the other hand, a higher threshold prioritizes precision, ensuring fewer incorrect alignments at the cost of reduced yield. The threshold can be adjusted based on the specific requirements of the aligned dataset and its intended downstream task. A detailed analysis of various DAC Thresholds is provided in Figure 7 and Table 4 in Appendix A.
@@ 5.1 Intrinsic Evaluation
Pralekha contains only parallel (alignable) documents, so aligning only these would not reflect a realistic and challenging scenario. To simulate real-world multilingual corpora setting, we sample random (unalignable) documents from Sangraha Unverified khan-etal-2024-indicllmsuite that cannot be aligned with those in Pralekha and inject them as noise into the dataset. We select a number of unalignable documents equal to 50% of the documents in Pralekha, resulting in a 1:2 ratio of unalignable to alignable documents during alignment. Table 1 reports the counts of unique alignable and unalignable documents for each language used in our experiments. Additionally, we compare the best-performing baseline with our proposed DAC approach on the CCAligned dataset el-kishky-etal-2020-ccaligned, using a similar data setting.
@@
We evaluate alignment at four granularities: sentence-level ($G=1$), chunk-level ($G=2,4,8$), and document-level ($G=|D|$). For $G=|D|$, we benchmark all pooling-based baselines described in Section 4.1. Unit embeddings are obtained using LaBSE [Uncaptioned image] feng-etal-2022-language and SONAR [Uncaptioned image] duquenne2023sonarsentencelevelmultimodallanguageagnostic. We build FAISS IndexFlatIP [^5] https://github.com/facebookresearch/faiss/wiki/The-index-factory indices and retrieve the top-$k$ nearest neighbors for each embedding with $k=16$. A DAC threshold of 0.1 is used to balance precision and recall, and a detailed threshold analysis is provided in Figure 7 and Table 4 in the Appendix.
@@
To assess the intrinsic performance of various document alignment methods, we evaluate three key metrics: precision, recall, and F1 score. While a high F1 score is generally desirable as it balances precision and recall, the ideal metric depends on the specific downstream task. We further measure computational time to analyze efficiency trade-offs across different granularities.
@@
5.2 Extrinsic Evaluation
Data Setting :
Implementation and Training :
We fine-tune two open-source LLMs: Llama-3.2-1B666huggingface.co/meta-llama/Llama-3.2-1B [Uncaptioned image] and Sarvam-1777huggingface.co/sarvamai/sarvam-1 [Uncaptioned image]. LLaMA-3.2-1B is a 1 billion parameter multilingual model released by Meta AI, while Sarvam-1 is a 2 billion parameter model developed by Sarvam AI, optimized for Indic languages. We used open-instruct [^8] https://github.com/allenai/open-instruct for full fine-tuning using the prompt-completion template. We used default hyperparameters recommended in open-instruct, most notably: learning rates of $5\times 10^{-5}$, maximum context (prompt+completion) lengths of 4096, Adam optimizer, and no weight decay. Training was performed on NVIDIA H100 GPUs with mixed-precision (bfloat16) support. We trained only bilingual models, a single model for English $\leftrightarrow$ X translation, for a total of 16 models (8 languages, 2 alignment approaches). Each model was trained for 1 epoch. We performed greedy decoding with a maximum of 4,096 generated tokens to maintain consistency across evaluations.